OpenStudy (anonymous):

The intensity of light at a distance r from a source is given by L = Ir^(−2,) where I is the illumination at the source. Starting with the values I = 20, r = 20, suppose we increase the distance by 4 and the illumination by 4. By (approximately) how much does the intensity of light change? dL= ???

OpenStudy (anonymous):

Given L(I,r)=Ir^(-2) deltaL=L(24,24)-L(20,20) deltaL=0.0416-0.05 = -0.0083

OpenStudy (anonymous):

i've done that but it says its incorrect !

OpenStudy (anonymous):

I'm not sure. It could be the rounding if you are doing it online.